# Steering Elongate Multi-legged Robots By Modulating Body Undulation Waves

### v1

@@ V Conclusion
The predicted turning arcs serve as fundamental motion primitives that are essential for a broad range of locomotion tasks, particularly those involving autonomous navigation in complex environments and obstacle avoidance. Because of the simple amplitude modulation scheme, implementation into closed-loop control involves the manipulation of only one variable for future real-time course correction. This ability to control planar trajectories serves as a substantial step forward in the development of practical, multi-legged robot systems for real world applications.
@@ ACKNOWLEDGMENT
We would like to thank Christopher Pierce for his technical consultation, Erik Teder for his consultation on figures and helpful discussion, and Juntao He for discussion. We would like to think the Institute for Robotics and Intelligent Machines for use of the College of Computing basement as a testing space. The authors received funding from NSF-Simons Southeast Center for Mathematics and Biology (Simons Foundation SFARI 594594), Army Research Office grant W911NF-11-1-0514, a Dunn Family Professorship, Ground Control Robotics, and a STTR Phase I (2335553) NSF grant.

### v2

@@ VI Conclusion and Future Work
This scheme builds upon previous work done in shape-based control of multi-legged robots that has proven to be robust across rugged terrain2[3]. By testing on a variety of terrains (carpet, hard floor, outdoors) and finding close agreement with theoretical predictions, we strengthened the locomotive assumptions in the model for moving in highly damped environments. Furthermore, the predicted turning arcs serve as fundamental motion primitives that are essential for a broad range of locomotion tasks, particularly those involving autonomous navigation in complex environments and obstacle avoidance. Because of the simple amplitude modulation scheme, implementation into closed-loop control involves the manipulation of only one variable for future real-time course correction. We successfully leveraged this fact for basic obstacle avoidance and wall-following algorithms and unlocked potential for work in full trajectory planning for this class of robots. This ability to control planar trajectories serves as a substantial step forward in the mobile applications of multi-legged robot systems, particularly in cluttered and difficult to navigate environments.
@@ ACKNOWLEDGMENT
We would like to thank Erik Teder for his consultation on figures and helpful discussion, and Juntao He for discussion. We would like to think the Institute for Robotics and Intelligent Machines for use of the College of Computing basement as a testing space. Additionally, we thank Ground Control Robotics LLC. for use of their robotic platform and support. The authors received funding from NSF-Simons Southeast Center for Mathematics and Biology (Simons Foundation SFARI 594594), Army Research Office grant W911NF-11-1-0514, a Dunn Family Professorship, Ground Control Robotics, and a STTR Phase I (2335553) NSF grant.